# SBC collaboration Low-Threshold Response of a Scintillating Xenon Bubble Chamber to Nuclear and Electronic Recoils

### v1

@@ IV.1 Nuclear Recoil Nucleation Efficiency
The measured neutron source and background single-bubble rates are shown as a function of in Fig. 3. The XeBC demonstrates an increased rate of bubbles over the background when exposed to 252Cf and the 88Y-Be photoneutron source, but no excess rate from the much weaker 207Bi-Be photoneutron source. The lack of evidence for nucleation by the weak 207Bi-Be source is to be expected — the original intent was to apply the scintillation veto to 207Bi-Be data to suppress the observed background, but as described above, the increased coincidence window precluded the use of the veto with a -ray source present. Bubble rates observed in background data are roughly constant over the range of explored here, while the rates of bubbles observed with 88Y-Be and 252Cf decrease as increases and the sensitivity to low-energy recoils diminishes. The maximum xenon recoil energy for NRs induced by 152 keV 88Y-Be neutrons is 4.6 keV. This gives an immediate upper bound on the actual bubble-nucleation threshold; because there is a rate above background, there must be some sensitivity to recoils below 4.6 keV.
To obtain a complete estimate of the NR nucleation efficiency as a function of recoil energy and thermodynamic state, a global analysis using all neutron calibration data gathered with the chamber is undertaken, closely following a similar analysis performed for the PICO dark matter experiment [18]. The data is binned by bubble count in each event (“multiplicity”), , and calibration source. The background data is binned in the same way, with a 1-to-1 map between source and background datasets. For consistency, the background and source data are required to be taken during the same xenon fill.
To extract a nucleation efficiency measurement, the expected energy spectrum and multiplicity of neutron interactions in the detector is required; this information is extracted from the detailed PoliMi output of the MCNPX-PoliMi simulations described in section III.3. In the limit where neutrons scatter at most once in the detector, the expected experimentally observed bubble rate can be calculated for a given bubble nucleation efficiency model as
@@
As no a priori functional form is known for the efficiency function , a generic model of piece-wise linear functions is used. Specifically, the efficiency function is modelled as a triangular mesh, with triangle vertices at fixed efficiencies (0, 0.2, 0.5, 0.8, and 1) and fixed (0.90 keV and 2.06 keV). Each rectangle in this grid is divided into two triangles by the diagonal from (low-, low-) to (high-, high-). The position of each vertex in is allowed to vary – giving 10 free parameters – but constrained such that the resulting efficiency function is monotonic in both and , and gives when . The vertex positions in efficiency are chosen to match Ref. [18], and the vertex positions in are chosen to match the highest/lowest in the data. Models with additional equally spaced vertices in are also considered. Choosing the optimal number of threshold () set-points is a balance between having a sufficiently flexible model while not over-fitting. To determine the ideal solution, the data is fit with models defined with up to five threshold set-points. The Akaike information criterion [35] is used to compare models with two or more threshold set-points, with two found to be the preferred number.
In addition to the 10 parameters of interest defining the nucleation efficiency model, nuisance parameters are included to reflect the neutron flux uncertainties from each source (9%, 13%, and 30% for the 252Cf, 88Y-Be, and 207Bi-Be sources, respectively) as well as a 10% common-mode uncertainty reflecting uncertainty on the mass of the visible xenon target. Each of these four nuisance parameters is implemented as a multiplier on the expected rates in the corresponding datasets.
@@
where () are the number of events observed with (without) the neutron source present in configuration , falling in bubble-multiplicity bin . The give the relative live-time of source-on to source-off data in each configuration. The floating parameters and indicate the efficiency curve and nuisance parameters, respectively, the latter each given a Gaussian prior with mean of zero and standard deviation one (giving a log-normal prior on the physical uncertainties they represent). The are the expected (i.e. simulated) number of bubble events generated by the source in configuration in bubble-multiplicity bin , given a particular efficiency curve and nuisance parameter set. Finally, the are implicit nuisance parameters for the background rate in each configuration and multiplicity bin. Given a hypothesis , the optimum is found analytically by solving . This leads to the non-negative, real solution
@@
Fitting the (10+4)-dimension model to the data is done with a custom Markov Chain Monte Carlo (MCMC) [36], using the algorithm described in Refs. [37, 18, 38] to determine both the best-fit and the boundary of the 1- region around the best fit. The nucleation efficiency model obtained from this analysis is shown in Fig. 4.
As found in previous measurements, the threshold for nucleation may deviate from . For the lower thermodynamic set-point, the best-fit threshold curve closely coincides with , the lower bound imposed on the model. The higher thermodynamic set-point curve is offset from by approximately 1 keV. A key result of this fit is that it indicates sensitivity to nuclear recoils of . Additionally, recent theoretical predictions of nucleation thresholds in various fluids using molecular dynamics simulations compare favourably with the present results [24].
A direct comparison of the best-fit model and each data point is given in Fig. 5. As expected, the 207Bi-Be data proves to be of little utility, with simulated count rates that are much smaller than measured backgrounds. Multi-bubble events in 88Y-Be data are also consistent with background at 2- or closer in all cases, meaning that constraints on the efficiency function are driven primarily by 88Y-Be single-bubble rates and 252Cf single- and multi-bubble rates. Post-fit constraints on nuisance parameters, given in Table 1, also show reasonable agreement with nominal values (all within 1.5 standard deviations).
To assess the quality of the fit result, a Monte Carlo study is carried as in Ref. [18]. Simulated datasets are generated using the best-fit model presented above, with random background and signal+background counts drawn for all experiments. 250 such datasets are generated, and then fit using the same procedure used for the real data. The resulting distribution of minimum for each fit is shown in Fig. 6, which is best fit by a distribution with 56 degrees of freedom. The distribution of of the MC datasets is then used to calculate a -value of for the best-fit model of the real data. This suggests that the generic model used in this analysis does adequately represent the data, and that the fit obtained is converged. Additional Monte Carlo studies to validate the model used in this work – in particular to demonstrate that this analysis paradigm produces unbiased results – were carried out in Ref. [38].
@@ IV.2 Gamma-Ray-Induced Bubble Nucleation
Raw bubble rates in the presence of the -ray sources are calculated in the same way as the rates for the nuclear recoil and background data described previously. Background-subtracted rates for -ray sources are calculated only for datasets that are accompanied by a background measurement under the same conditions, and only if both the -ray source and background datasets have greater than 10 bubbles. These requirements are met only for data at thresholds 0.90 keV, due to challenges acquiring significant live time at fixed threshold in this device at higher degrees of superheat. Data at thresholds from 0.50–0.90 keV are still used without background subtraction to give upper limits on -induced bubble nucleation rates.
@@
The strongest upper limits, of nucleation probability, are produced by 88Y data, due to the high activity source. At the lowest achievable Seitz threshold with this setup, 0.50 keV, an upper limit of is set using 88Y without background subtraction. Sensitivities for the background-subtracted measurements, calculated according to the Feldman-Cousins procedure [39] by simulating 1000 trials per experiment and measuring the average upper limit attained in the simulated trials, are consistent with the observed limits.
These results may be compared to the PICO Auger-cascade bubble nucleation model [17] for xenon-contaminated C3F8 at an operating pressure of 1.7 bara, which provides a prediction for the nucleation rate per xenon photoabsorption as a function of , shown in Fig. 7. The -ray rejection in xenon is found to be at least six orders of magnitude stronger than the projected value from xenon-contaminated C3F8 for thresholds around 1 keV.
@@ V Conclusion and Discussion
At the same time, electronic recoils continue not to nucleate bubbles in xenon at down to at least 0.5 keV, in agreement with past literature on liquid xenon and in extreme tension with electronic recoil bubble nucleation in C3F8. The authors suggest an interpretation of this phenomenon similar to that found in [40] — namely that bubble nucleation by -electrons and Auger cascades requires the presence of molecular degrees of freedom, which provide a channel for electrons to directly heat the fluid, e.g. by exciting molecular vibrational states. In the absence of these degrees of freedom, electrons can lose energy only through ionization and radiation, including through the scintillation emission noble liquids are most known for.
@@ VI Acknowledgments
We are grateful to the Northwestern University department of Physics for their support in running this detector. This work was supported by DOE Office of Science grants DE-SC0015910, DE-SC0017815, DE-SC0024254, and DE-SC0011702, National Science Foundation grants DMR-1936432, PHY-2310112, and PHY-2411655, Project No. CONACYT CB-2017-2018/A1-S-8960, DGAPA UNAM Grant No. PAPIIT IN105923, and Fundación Marcos Moshinsky. We acknowledge support from the Canada First Research Excellence Fund through the Arthur B. McDonald Canadian Astroparticle Physics Research Institute and the Natural Sciences and Engineering Research Council of Canada (NSERC).
We also thank the Digital Research Alliance of Canada [43] and the Centre for Advanced Computing, ACENET, Calcul Québec, Compute Ontario, and WestGrid for computational support. The work of M. Bressler was supported by the Department of Energy Office of Science Graduate Instrumentation Research Award (GIRA). The work of D. Durnford and A. de St. Croix was additionally supported by the NSERC Canada Graduate Scholarships – Doctoral program (CGSD).

### v2

@@ IV.1 Nuclear recoil nucleation efficiency
The measured neutron source and background single-bubble rates are shown as a function of in Fig. 3. The XeBC demonstrates an increased rate of bubbles over the background when exposed to 252Cf and the 88Y-Be photoneutron source, but no excess rate from the much weaker 207Bi-Be photoneutron source. The lack of evidence for nucleation by the weak 207Bi-Be source is to be expected – the original intent was to apply the scintillation veto to 207Bi-Be data to suppress the observed background, but as described above, the increased coincidence window precluded the use of the veto with a -ray source present. Bubble rates observed in background data are roughly constant over the range of explored here, while the rates of bubbles observed with 88Y-Be and 252Cf decrease as increases and the sensitivity to low-energy recoils diminishes. The maximum xenon recoil energy for NRs induced by 152 keV 88Y-Be neutrons is 4.6 keV. This gives an immediate upper bound on the actual bubble-nucleation threshold; because there is a rate above background, there must be some sensitivity to recoils below 4.6 keV.
Figure 3 illustrates nucleation rates versus the Seitz threshold for fixed nuclear recoil spectra, but physics results from dark matter searches or CENS measurements depend on knowing bubble nucleation efficiency as a function of recoil energy at a fixed thermodynamic state. Some past efforts at bubble chamber calibration have performed this conversion by assuming specific functional forms for nucleation efficiency that link dependence on recoil energy to dependence on the Seitz threshold [7, 27], but the physical basis for such assumptions is unclear. While ongoing theoretical efforts attempt to better constrain nucleation efficiency models using a combination of recoil cascade simulations and molecular dynamics evolutions [28, 29], recent experimental efforts have shifted toward purely data-driven constraints on nucleation efficiency that gain their constraining power from measurements of multiple recoil spectra, rather than measurements taken at multiple thermodynamic states. The analysis presented here takes the approach developed in [22], making the same minimal set of assumptions: thermodynamic states with the same are assumed to give the same nuclear recoil response (an assumption supported by data in [22] for the range of pressures considered in this work); efficiency is assumed to increase as recoil energy increases or decreases; and efficiency is assumed to be zero when recoil energy is less than .
To obtain a complete estimate of the NR nucleation efficiency as a function of recoil energy and , a global analysis using all neutron calibration data gathered with the chamber is undertaken, closely following the analysis performed in Ref. [22]. The data is binned by bubble count in each event (“multiplicity”), , and calibration source. The background data is binned in the same way, with a 1-to-1 map between source and background datasets. For consistency, the background and source data are required to be taken during the same xenon fill.
To extract a nucleation efficiency measurement, the expected energy spectrum and multiplicity of neutron interactions in the detector is required; this information is extracted from the detailed PoliMi output of the MCNPX-PoliMi simulations described in Sec. III.3. In the limit where neutrons scatter at most once in the detector, the expected experimentally observed bubble rate can be calculated for a given bubble nucleation efficiency model as
@@
As no a priori functional form is known for the efficiency function , a generic model of piecewise linear functions is used. Specifically, the efficiency function is modeled as a triangular mesh, with triangle vertices at fixed efficiencies (0, 0.2, 0.5, 0.8, and 1) and fixed (0.90 keV and 2.06 keV). Each rectangle in this grid is divided into two triangles by the diagonal from (low-, low-) to (high-, high-). The position of each vertex in is allowed to vary – giving 10 free parameters – but constrained such that the resulting efficiency function is monotonic in both and , and gives when . The vertex positions in efficiency are chosen to match Ref. [22], and the vertex positions in are chosen to match the highest/lowest in the data. Models with additional equally spaced vertices in are also considered. Choosing the optimal number of threshold () set-points is a balance between having a sufficiently flexible model while not overfitting. To determine the ideal solution, the data is fit with models defined with up to five threshold set-points. The Akaike information criterion [39] is used to compare models with two or more threshold set-points, with two found to be the preferred number.
In addition to the ten parameters of interest defining the nucleation efficiency model, nuisance parameters are included to reflect the neutron flux uncertainties from each source (9%, 13%, and 30% for the 252Cf, 88Y-Be, and 207Bi-Be sources, respectively) as well as a 10% common-mode uncertainty reflecting uncertainty on the mass of the visible xenon target. Each of these four nuisance parameters is implemented as a multiplier on the expected rates in the corresponding datasets.
@@
where () are the number of events observed with (without) the neutron source present in configuration , falling in bubble-multiplicity bin . The give the relative live time of source-on to source-off data in each configuration. The floating parameters and indicate the efficiency curve and nuisance parameters, respectively, the latter each given a Gaussian prior with mean of zero and standard deviation one (giving a log-normal prior on the physical uncertainties they represent). The are the expected (i.e., simulated) number of bubble events generated by the source in configuration in bubble-multiplicity bin , given a particular efficiency curve and nuisance parameter set. Finally, the are implicit nuisance parameters for the background rate in each configuration and multiplicity bin. Given a hypothesis , the optimum is found analytically by solving . This leads to the non-negative, real solution
@@
Fitting the (10+4)-dimension model to the data is done with a custom Markov chain Monte Carlo (MCMC) [40], using the algorithm described in Refs. [41, 22, 16] to determine both the best fit and the boundary of the 1- region around the best fit. The nucleation efficiency model obtained from this analysis is shown in Fig. 4.
As found in previous measurements, the threshold for nucleation may deviate from . For the lower thermodynamic set-point, the best-fit threshold curve closely coincides with , the lower bound imposed on the model. The higher thermodynamic set-point curve is offset from by approximately 1 keV. A key result of this fit is that it indicates sensitivity to nuclear recoils of . Additionally, recent theoretical predictions of nucleation thresholds in various fluids using molecular dynamics simulations compare favorably with the present results [28].
A direct comparison of the best-fit model and each data point is given in Fig. 5. As expected, the 207Bi-Be data proves to be of little utility, with simulated count rates that are much smaller than measured backgrounds. Multi-bubble events in 88Y-Be data are also consistent with background at 2- or closer in all cases, meaning that constraints on the efficiency function are driven primarily by 88Y-Be single-bubble rates and 252Cf single- and multi-bubble rates. Postfit constraints on nuisance parameters, given in Table 1, also show reasonable agreement with nominal values (all within 1.5 standard deviations).
To assess the quality of the fit result, a Monte Carlo study is carried as in Ref. [22]. Simulated datasets are generated using the best-fit model presented above, with random background and signal+background counts drawn for all experiments. There are 250 such datasets generated, and then fit using the same procedure used for the real data. The resulting distribution of minimum for each fit is shown in Fig. 6, which is best fit by a distribution with 56 degrees of freedom. The distribution of of the MC datasets is then used to calculate a -value of for the best-fit model of the real data. This suggests that the generic model used in this analysis does adequately represent the data, and that the fit obtained is converged. Additional Monte Carlo studies to validate the model used in this work – in particular to demonstrate that this analysis paradigm produces unbiased results – were carried out in Ref. [16].
@@ IV.2 Gamma-ray-induced bubble nucleation
Raw bubble rates in the presence of the -ray sources are calculated in the same way as the rates for the nuclear recoil and background data described previously. Background-subtracted rates for -ray sources are calculated only for datasets that are accompanied by a background measurement under the same conditions, and only if both the -ray source and background datasets have greater than ten bubbles. These requirements are met only for data at thresholds 0.90 keV, due to challenges acquiring significant live time at the fixed threshold in this device at higher degrees of superheat. Data at thresholds from 0.50–0.90 keV are still used without background subtraction to give upper limits on -induced bubble nucleation rates.
@@
The strongest upper limits – of nucleation probability – are produced by 88Y data, due to the high activity of the source. At the lowest achievable Seitz threshold with this setup, 0.50 keV, an upper limit of is set using 88Y without background subtraction. Sensitivities for the background-subtracted measurements, calculated according to the Feldman-Cousins procedure [42] by simulating 1000 trials per experiment and measuring the average upper limit attained in the simulated trials, are consistent with the observed limits.
These results may be compared to the PICO Auger-cascade bubble nucleation model [21] for xenon-contaminated C3F8 at an operating pressure of 1.7 bara, which provides a prediction for the nucleation rate per xenon photoabsorption as a function of , shown in Fig. 7. The -ray rejection in xenon is found to be at least 6 orders of magnitude stronger than the projected value from xenon-contaminated C3F8 for thresholds around 1 keV.
@@ V Conclusion and Discussion
At the same time, electronic recoils continue not to nucleate bubbles in xenon at down to at least 0.5 keV, in agreement with past literature on liquid xenon and in extreme tension with electronic recoil bubble nucleation in C3F8. The authors suggest an interpretation of this phenomenon similar to that found in [43] – namely that bubble nucleation by -electrons and Auger cascades requires the presence of molecular degrees of freedom, which provide a channel for electrons to directly heat the fluid, e.g. by exciting molecular vibrational states. In the absence of these degrees of freedom, electrons can lose energy only through ionization and radiation, including through the scintillation emission noble liquids are most known for. A key open question is whether this insensitivity to electron recoils persists down to the homogeneous nucleation limit, the point at which random fluctuations in density begin to nucleate bubbles in the superheated liquid at a measurable rate. This limit is reached at Seitz thresholds of 75 eV and 40 eV in superheated xenon and argon, respectively [44].
@@ VI Acknowledgments
We are grateful to the Northwestern University department of Physics and Astronomy and to a number of students including Theo Baker, Jon Chen, Trent Cwiok, Allison Grimsted, Jared Gupta, Miaotianzi Jin, Jacob McLaughlin, Ricky Puig, Dylan Temples, and David Velasco for their support in running this detector. This work was supported by DOE Office of Science Grants No. DE-SC0015910, No. E-SC0017815, No. DE-SC0024254, and Do. DE-SC0011702, National Science Foundation Grants No. DMR-1936432, No. PHY-2310112, and No. PHY-2411655, Project No. CONACYT CB-2017-2018/A1-S-8960, DGAPA UNAM Grant No. PAPIIT IN105923, and Fundación Marcos Moshinsky. We acknowledge support from the Canada First Research Excellence Fund through the Arthur B. McDonald Canadian Astroparticle Physics Research Institute and the Natural Sciences and Engineering Research Council of Canada (NSERC).
We also thank the Digital Research Alliance of Canada [45] and the Centre for Advanced Computing, ACENET, Calcul Québec, Compute Ontario, and WestGrid for computational support. The work of M. Bressler was supported by the Department of Energy Office of Science Graduate Instrumentation Research Award. The work of D. Durnford and A. de St. Croix was additionally supported by the NSERC Canada Graduate Scholarships – Doctoral program.